# Empowering Molecule Discovery for Molecule-Caption Translation with Large Language Models: A ChatGPT Perspective

### v1

@@ Abstract
Molecule discovery plays a crucial role in various scientific fields, advancing the design of tailored materials and drugs. Traditional methods for molecule discovery follow a trial-and-error process, which are both time-consuming and costly, while computational approaches such as artificial intelligence (AI) have emerged as revolutionary tools to expedite various tasks, like molecule-caption translation. Despite the importance of molecule-caption translation for molecule discovery, most of the existing methods heavily rely on domain experts, require excessive computational cost, and suffer from poor performance. On the other hand, Large Language Models (LLMs), like ChatGPT, have shown remarkable performance in various cross-modal tasks due to their great powerful capabilities in natural language understanding, generalization, and reasoning, which provides unprecedented opportunities to advance molecule discovery. To address the above limitations, in this work, we propose a novel LLMs-based framework (MolReGPT) for molecule-caption translation, where a retrieval-based prompt paradigm is introduced to empower molecule discovery with LLMs like ChatGPT without fine-tuning. More specifically, MolReGPT leverages the principle of molecular similarity to retrieve similar molecules and their text descriptions from a local database to ground the generation of LLMs through in-context few-shot molecule learning. We evaluate the effectiveness of MolReGPT via molecule-caption translation, which includes molecule understanding and text-based molecule generation. Experimental results show that MolReGPT outperforms fine-tuned models like MolT5-base without any additional training. To the best of our knowledge, MolReGPT is the first work to leverage LLMs in molecule-caption translation for advancing molecule discovery. Our implementation is available at: https://github.com/phenixace/MolReGPT
Machine Learning, ICML
@@ 1 Introduction
Molecules are the fundamental building blocks of matter, comprising the intricate fabric of the world around us. As the foundation of all chemical compounds, molecules are composed of two or more atoms that are chemically bonded together, and they retain the unique chemical properties dictated by their specific structures (Xu et al., 2023). With a comprehensive understanding of molecules, scientists can effectively design materials, drugs, and products with tailored characteristics and functionalities, impacting a variety of crucial fields such as chemistry (Wang et al., 2023; Cuzzucoli Crucitti et al., 2023; Weng et al., 2021), pharmacology (Patani & LaVoie, 1996; Anderson, 2003; Ding et al., 2019), material science (Curtarolo et al., 2013; Higuchi et al., 2023), and environmental science (Ali et al., 2023; Lv et al., 2023). One notable example is in the pharmaceutical industry during the COVID-19 pandemic, where the discovery of new molecules has the potential to revolutionize not only the development of groundbreaking treatments, therapies (Gupta et al., 2023), and vaccines against viruses but also a wide range of other diseases in the coming decade (Osamor et al., 2023).
However, traditional molecule discovery lies in the long, expensive, and failure-prone process that requires navigating a complex landscape of molecule structures and biological interactions, with limitations in scalability, precision, and data management (Hajduk & Greer, 2007). To overcome these challenges, computational technologies such as artificial intelligence (AI) have emerged as powerful tools to expedite the discovery of new molecules (Urbina & Ekins, 2022). Specifically, molecules can be represented as simplified molecular-input line-entry system (SMILES) strings (Weininger, 1988; Cao et al., 2022). As shown in Figure 1 (a), the structure of Phenol can be represented as a SMILES sequence, which is made of a Benzene ring and a Hydroxy. Such SMILES representations can be effectively processed by deep sequence models like Recurrent Neural Networks (Arús-Pous et al., 2019; Grisoni et al., 2020) and Transformers (Honda et al., 2019; Yoshikai et al., 2023). These AI-powered models enable researchers to understand molecular properties and functionalities and create promising compounds in a more efficient and cost-effective manner. For example, in order to generate new molecules and better understand them, a novel task that translates between molecules and natural language has been proposed by using language models like Text2Mol  (Edwards et al., 2021) and MolT5  (Edwards et al., 2022). It consists of two sub-tasks: molecule captioning (Mol2Cap) and text-based molecule generation (Cap2Mol). More specifically, as shown in Figure 1 (b-c), the goal of molecule captioning is to generate a text caption to describe a SMILES string of the molecule for providing humans with a better understanding of molecule, while text-based molecule generation aims to generate the corresponding molecule (i.e., SMILES string) based on a given natural language description (e.g., properties and functional groups). Despite the impressive progress that has been made in the molecule-caption translation task, the majority of existing advanced approaches suffer from several limitations (Edwards et al., 2021, 2022; Su et al., 2022). First, the design of such model architecture in molecule-caption translation heavily relies on domain experts, which can significantly limit the development and deployment of AI-powered molecule discovery. Second, most existing methods follow the “pre-train&fine-tuning” paradigm for molecule-caption translation, which requires excessive computational costs. Third, existing approaches such as Text2Mol and MolT5 fall short in their inability to reason on complex tasks and generalize to unseen examples. Therefore, it is desired to design a novel paradigm for molecule-caption translation in molecule discovery.
Refer to caption (a) Molecule Representations. Refer to caption (b) Molecule Captioning. Refer to caption (c) Text-based Molecule Generation.
@@
Figure 1: An illustration on translation between molecule and natural language in molecule discovery. (a) A molecule can be denoted as a chemical formula, SMILES string, and 2D molecule graph. (b) Molecule captioning aims to generate a text caption to describe a molecule for humans’ better understanding. (c) Given a text description, text-based molecule generation is used to generate a corresponding molecule. (d) Large language models (e.g., ChatGPT) can perform molecule captioning and text-based molecule generation with corresponding well-designed prompts.
@@
Recently, Large Language Models (LLMs), scaling up their weights to the billion level, have achieved tremendous success not only in the field of Natural Language Processing (NLP) but also in some cross-modal areas like computer vision (Zhu et al., 2023), recommender systems (Bao et al., 2023), and molecule discovery (Edwards et al., 2022). For example, ChemGPT (Frey et al., 2022), a variant of GPT model with more than one billion parameters, is introduced to understand and generate small molecules in chemistry. Meanwhile, in addition to the impressive capabilities in natural language understanding and generation, LLMs also demonstrate their powerful generalization and reasoning capabilities (Rubin et al., 2022; Min et al., 2022), which can generalize to other unseen tasks by specific task context (In-Context Learning, ICL) without being fine-tuned and largely reduce computational cost. Therefore, LLMs provide unprecedented potential to advance molecule discovery, specifically the task of molecule-caption translation.
@@
Although building specific LLMs in molecule discovery has immense potential for advancing scientific research, we also face significant challenges. First, due to privacy and security concerns, many advanced large language models (e.g., ChatGPT and GPT4.0) are not publicly available, where LLMs’ architectures and parameters are not released publicly for fine-tuning in downstream tasks. Second, owing to their complex architectures and the extensive data required, training advanced LLMs requires significant computing resources, leading to high costs and substantial energy consumption. For instance, it has been reported that the cost of one single training session for GPT-3 exceeds 1 million. As a result, it is very challenging for us to re-design our own LLMs with pre-training and fine-tuning in specific downstream tasks. At last, it is crucial to design proper guidelines/prompts with high-quality few-shot examples to improve LLMs’ generalization and reasoning capabilities for molecule discovery.
@@
To address such challenges, as the early exploration attempt to take advantage of the powerful capabilities of LLMs in the molecule discovery field, in this work, we propose a novel solution to teach LLMs with prompts for translating between molecules and natural language, as illustrated in Figure 1 (d). More specifically, inspired by the latest ChatGPT, a retrieval-based prompt paradigm through in-context learning (ICL) is developed to conduct two sub-tasks (i.e., molecule captioning and text-based molecule generation) without fine-tuning the LLMs, where n-shot examples are retrieved to augment the prompt instances via BM25-based caption ranking and Morgan Fingerprints-based molecule ranking. Experiments show that MolReGPT can achieve Text2Mol scores of 0.560 in Mol2Cap generation and 0.571 in Cap2Mol generation, which surpasses the fine-tuned MolT5-base in both sub-tasks of molecule-caption translation. Notably, MolReGPT even outperforms MolT5 largely in text-based molecule generation, increasing the Text2Mol metric by 3%. Note that all of these improvements from our proposed method are achieved without any fine-tuning steps.
@@
  • We introduce a principle strategy based on LLMs to perform translation between molecules and natural language for molecule discovery. To the best of our knowledge, we are the first to investigate molecule-caption translation by employing LLMs.
  • We develop a novel framework (MolReGPT) to empower LLMs like ChatGPT to perform molecule captioning and text-based molecule generation without being fine-tuned, where a retrieval-based prompt paradigm through in-context learning is developed to explicitly guide the generation process.
  • Comprehensive experiments on a real-world dataset demonstrate the effectiveness of the proposed method on molecule captioning and text-based molecule generation tasks, surpassing even fine-tuned models such as T5-base and MolT5-base.
@@ 2 Related Work
In this section, we briefly review related work about molecule-caption translation tasks in molecule discovery as well as the advanced LLMs techniques.
2.1 Molecule Discovery
Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design (Du et al., 2022). In recent decades, AI-powered approaches have emerged as mainstream techniques to revolutionize the process of molecule discovery (Hu et al., 2023; Fan et al., 2023). For instance, SMILES-based Variational Autoencoders (VAEs) methods such as ChemVAE (Gómez-Bombarelli et al., 2018), SD-VAE (Dai et al., 2018), and GrammarVAE (Kusner et al., 2017), employ a VAE-based model that encodes and decodes SMILES strings, generating new molecules (strings) by decoding from a Gaussian prior. In terms of molecular string representation, existing studies have explored advanced deep representation methods from other fields, including Convolutional Neural Network (CNN) (Peng & Zhao, 2019; Le et al., 2019), Recurrent Neural Network (RNN) (Grisoni et al., 2020; Amabilino et al., 2020), and Transformer (Bagal et al., 2021; Wang et al., 2021).
@@ 2.1 Molecule Discovery
More recently, as a new task in molecule discovery, Text2Mol (Edwards et al., 2021) is introduced to retrieve molecules using natural language descriptions as search queries, in which a paired dataset of molecules and their corresponding text descriptions are constructed, enabling the learning of a shared semantic embedding space for retrieval. KV-PLM (Zeng et al., 2022) develops a knowledgeable machine reading system pre-trained on a domain corpus, in which SMILES strings are inserted and link molecule structures with biomedical text. What’s more, a self-supervised learning framework MolT5 (Edwards et al., 2022) is proposed to pre-train on a substantial volume of unlabeled language text and SMILES strings, enhancing the molecule-caption translation task, such as molecule captioning and text-based molecule generation. MoMu (Su et al., 2022) bridges molecular graphs and natural language by pre-training molecular graphs and their semantically related text data through comparative learning.
2.2 Large Language Models
@@
Large Language models (LLMs) have been a trending topic in recent years, with numerous studies exploring their capabilities and potential applications. One of the most well-known LLMs is the GPT family (Radford et al., 2018, 2019; Brown et al., 2020; Ouyang et al., 2022), which has played a pivotal role in advancing the field of generative language models. As a representative of the GPT family, ChatGPT is specifically fine-tuned for conversational purposes, which can generate impressively human-like responses (Leiter et al., 2023). In addition, other LLMs, such as LaMDA (Thoppilan et al., 2022), PaLM (Chowdhery et al., 2022), and Vicuna (Chiang et al., 2023), also show a decent performance.
The power of LLMs is far beyond language generation but also lies in their ability to learn from context, namely the ability of in-context learning. Several works have explored the utilization of in-context learning from various tasks, such as KATE (Liu et al., 2021) and AutoCoT (Zhang et al., 2022). These works show that through in-context learning, LLMs can adapt to new tasks based on the context provided in the input, eliminating the need for explicit fine-tuning on specific tasks.
@@
In addition to NLP tasks, LLMs have also shown remarkable potential in various molecule discovery tasks, such as molecule understanding (Bran et al., 2023; White, 2023). For instance, ChemBERTa (Chithrananda et al., 2020) leverages pre-training on an extensive corpus of chemical texts, enabling it to comprehend the structure and properties of chemical compounds. Another notable example is MoleculeSTM (Liu et al., 2022), which employs in-context learning in conjunction with LLMs. This approach facilitates a deeper understanding of the relationships between chemical structures and their corresponding textual descriptions. Furthermore, ChemGPT (Frey et al., 2022) represents a variant of the GPT model specifically trained on chemical data. Through the application of in-context learning, ChemGPT is capable of generating novel chemical structures and accurately predicting their properties. MolT5 (Edwards et al., 2022) shows that LLMs can perform the cross-modal transition task between molecule and text (i.e. molecule captioning task and text-based molecule generation task), which is one of the most closely related attempts to ours. Note that MolT5 needs to pre-train and fine-tune LLMs for translating between molecules and natural language, leading to huge computational costs. In this paper, we propose a novel framework to empower LLMs like ChatGPT to perform molecule captioning and text-based molecule generation without being fine-tuned.
@@ 3 MolReGPT
In this section, we aim to introduce our proposed method (MolReGPT) as a novel solution to empower molecule discovery for molecule-caption translation with LLMs like ChatGPT. We will first introduce an overview of the proposed framework, and then detail each model component.
Figure 2: This diagram shows the workflow of MolReGPT. MolReGPT consists of four main stages. In stage 1, Molecule-Caption Retrieval is employed to find $n$ best-matched examples from the local database. Then in stage 2, Prompt Management helps construct the system prompt with the retrieved molecule-caption pairs. Following this, LLMs perform In-Context Few-Shot Molecule Learning based on the provided system prompt and user input prompt. Finally, Generation Calibration is conducted to ensure desired output.
3.1 An Overview
Due to the huge computation costs, training or fine-tuning LLMs on the domain-specific corpus from the molecule discovery field is often infeasible in practice. To address such limitations, we investigate leveraging the great capabilities of LLMs without changing the LLMs, where we propose a novel framework (MolReGPT) to empower ChatGPT with the ability of molecule-caption translation for molecule discovery. More specifically, in order to improve the quality of guidelines/prompts, a retrieval-based prompt paradigm under in-context learning is introduced to teach ChatGPT to conduct two molecule-related tasks: molecule captioning (Mol2Cap) and text-based molecule generation (Cap2Mol). The framework of MolReGPT is shown in Figure 2, consisting of four main stages: Molecule-Caption Retrieval, Prompt Management, In-Context Few-Shot Molecule Learning, and Generation Calibration, following the workflow of pre-processing, querying, and post-processing. The first stage, Molecule-Caption Retrieval, is used to retrieve the $n$ most similar examples (i.e., few-shot examples) from human-annotated datasets (i.e., molecule-caption pairs database) for augmenting the prompt instances. The second stage is Prompt Management, which is executed to construct the system prompt as evidence for successive in-context learning. After that, both the system and the user input prompt are sent to query LLMs such as ChatGPT to perform In-Context Few-Shot Molecule Learning without fine-tuning LLMs for the molecule-caption translation task in the molecule discovery field. Valid responses are expected in the pre-defined JSON format, while there may be instances where the language models generate unexpected outputs. The last stage is Generation Calibration, which is deployed to calibrate the original responses for the validity of the outputs.

### v2

@@ Abstract
[…]
Index Terms:
Drug Discovery, Large Language Models (LLMs), In-context Learning, Retrieval Augmented Generation.
@@ 1 Introduction
(a) Molecule Representations.
(b) Molecule Captioning.
(c) Text-based Molecule Generation.
@@
Figure 1: An illustration of molecule-caption translation. (a) Representations of a molecule. (b) Molecule captioning (Mol2Cap) aims to generate a text caption describing the features of the molecule. (c) Text-based molecule generation (Cap2Mol) generates a corresponding molecule to the given caption. (d) Large language models (e.g., ChatGPT) can perform Mol2Cap and Cap2Mol with well-designed prompts.
As the foundation of chemical compounds, molecules are composed of two or more atoms that are chemically bonded together, denoting their unique chemical properties dictated by their specific structures [1]. With a comprehensive understanding of molecules, scientists can effectively design materials, drugs, and products with tailored characteristics and functionalities, impacting a variety of crucial fields such as chemistry [2], pharmacology [3], and material science [4].
Recently, computational technologies such as artificial intelligence (AI) have emerged as powerful tools to expedite the discovery of new molecules [5]. Specifically, molecules can be represented as simplified molecular-input line-entry system (SMILES) strings [6], illustrated in Figure 1 (a), which can be effectively processed by deep sequence models like Recurrent Neural Networks [7] and Transformers [8]. These AI-powered models enable researchers to understand molecular properties and functionalities and thus create promising compounds in a more efficient and cost-effective manner. For example, in order to generate new molecules and better comprehend them, a novel task that translates between molecules and natural language has been proposed by using language models like Text2Mol  [9] and MolT5  [10]. It consists of two sub-tasks: molecule captioning (Mol2Cap) and text-based molecule generation (Cap2Mol). As shown in Figure 1 (b-c), the goal of Mol2Cap is to generate a text caption describing molecule features (e.g. structures and characteristics). Specifically, the text caption first demonstrates the molecule structure by describing functional group positions and the according IUPAC name. Then, the characteristics of the molecule will be discussed, including the family that it belongs to and the chemical features that affect its practical use. On the other hand, Cap2Mol aims to generate the corresponding molecule (i.e., SMILES string) based on the given text caption, where the structure and feature information of the molecules could be used to infer the SMILES representation of the molecule. For example, given the molecule ”CCCCCCCCCCCCCCCCCCCCCCCCCCCCC”, its caption writes ”The molecule is a straight-chain alkane comprising of 29 carbon atoms. It has a role as a plant metabolite and a volatile oil component”, where ”straight-chain alkane” and ”29 carbon atoms” directly describe the molecule structure, while the last sentence shows the functions of the molecule.
Despite the impressive progress that has been made in the molecule-caption translation task [9, 10, 11], the majority of existing advanced approaches still suffer from several limitations. First, the design of model architectures heavily relies on the labour of domain experts, which can significantly limit the development of AI-powered molecule discovery. Second, most existing methods follow the pre-train&fine-tuning paradigm for molecule-caption translation, putting demanding requirements on computational and domain resources. Third, existing approaches such as Text2Mol and MolT5 fall short in their inability to generalize to unseen examples. Therefore, it is desired to design a novel paradigm for molecule-caption translation.
@@
Recently, Large Language Models (LLMs), scaling up their weights to the billion level, have achieved tremendous success not only in the field of Natural Language Processing (NLP) but also in some cross-modal areas like computer vision [12], recommender systems [13], and molecule discovery [10]. Meanwhile, in addition to the impressive capabilities in natural language understanding and generation, LLMs also demonstrate their powerful generalization and reasoning capabilities [14, 15], which can generalize to other unseen tasks via in-context learning without the necessity of being fine-tuned, largely reducing computational cost. Therefore, LLMs provide unprecedented potential to advance molecule discovery, specifically the task of molecule-caption translation.
@@
Although building specific LLMs in molecule discovery has immense potential for advancing scientific research, we also face significant challenges. First, due to privacy and security concerns, many advanced large language models (e.g., GPT-3.5 and GPT-4) are not publicly available, where LLMs’ architectures and parameters are not released publicly for fine-tuning in downstream tasks. Second, owing to their complex architectures and the extensive data required, training advanced LLMs requires significant computing resources and domain corpus, leading to high costs and substantial energy consumption. For instance, it has been reported that the cost of one single training session for GPT-3 exceeds 1 million US dollars. As a result, it is very challenging for us to re-design our own LLMs with pre-training and fine-tuning in specific downstream tasks. At last, it still lacks proper guidelines/paradigms for scientific researchers to make use of powerful LLMs like ChatGPT to enhance their own study in molecule discovery.
@@
To address such challenges, as the early exploration attempt to take advantage of the powerful capabilities of LLMs in the molecule discovery field, in this work, we propose a novel solution to teach LLMs with specific in-context examples for translating between molecules and natural language, as illustrated in Figure 1 (d). More specifically, inspired by the latest ChatGPT, a retrieval-based In-Context Few-Shot Molecule Learning paradigm is developed to conduct two sub-tasks (i.e., Mol2Cap and Cap2Mol) without fine-tuning the LLMs, where n similar molecule-caption pairs are retrieved as context instances under the guidance of molecular similarity, including BM25-based caption retrieval and Morgan Fingerprints-based molecule retrieval. Experiments show that MolReGPT (GPT-4-0314) achieves Text2Mol scores of 0.585 in Mol2Cap and 0.593 in Cap2Mol, showing comparable performance to MolT5-large in Mol2Cap and even outperforming MolT5-large in Cap2Mol without any fine-tuning steps, increasing the Text2Mol metric by 0.5% and 6%, respectively.
@@
  • We introduce a principle strategy based on LLMs, In-Context Few-Shot Molecule Learning, to perform translation between molecules and natural language for molecule discovery. To the best of our knowledge, we are the first to employ the in-context learning ability of LLMs in molecule-caption translation. Our work expands the application scope of LLMs and provides valuable insights into how these LLMs can be adapted for specific scientific tasks.
  • We develop a novel framework (MolReGPT) to empower LLMs like ChatGPT for scientific purposes without being pre-trained or fine-tuned on domain-specific corpora. By enabling LLMs to understand and generate meaningful molecular descriptions and molecule structures, we pave the way for AI-assisted drug discovery and design. MolReGPT has the potential to accelerate the development of new pharmaceuticals and improve the efficiency of molecular research.
  • We conduct comprehensive experiments on a real-world dataset with molecule-caption pairs to demonstrate the effectiveness and mechanisms of the proposed method on Mol2Cap and Cap2Mol tasks. The results show that our method could enable GPT4-0314 to achieve comparable performance to MolT5-large.
@@ 2 Related Work
Molecule Discovery. In recent decades, AI-powered approaches have emerged as mainstream techniques to revolutionize the process of molecule discovery [16, 17]. existing studies have explored advanced deep representation methods from other fields, including Convolutional Neural Network (CNN) [18, 19], Recurrent Neural Network (RNN) [20, 21], and Transformer [22, 23]. More recently, as a new task in molecule discovery, Text2Mol [9] is introduced to retrieve molecules using natural language descriptions as search queries, in which a paired dataset of molecules and their corresponding text descriptions are constructed, enabling the learning of a shared semantic embedding space for retrieval. KV-PLM [24] develops a knowledgeable machine reading system pre-trained on a domain corpus, in which SMILES strings are inserted and link molecule structures with biomedical text. What’s more, a self-supervised learning framework MolT5 [10] is proposed to pre-train on a substantial volume of unlabeled language text and SMILES strings, enhancing the molecule-caption translation task, such as molecule captioning and text-based molecule generation. MoMu [11] bridges molecular graphs and natural language by pre-training molecular graphs and their semantically related text data through comparative learning.
@@
Large Language Models (LLMs). LLMs have been a trending topic in recent years, with numerous studies exploring their capabilities and potential applications. One of the most well-known LLMs is the GPT family [25, 26, 27], which has played a critical role in advancing the field of generative language models. As a representative of the GPT family, ChatGPT is specifically fine-tuned for conversational purposes, which can generate impressively human-like responses [28]. In addition, other LLMs, such as LaMDA [29], PaLM [30], and Vicuna [31], also show a decent performance.
@@
In addition to NLP tasks, LLMs have also shown remarkable potential in various molecule discovery tasks, such as molecule understanding [32, 33]. For instance, ChemBERTa [34] leverages pre-training on an extensive corpus of chemical texts, enabling it to comprehend the structure and properties of chemical compounds. Another notable example is MoleculeSTM [35], which employs in-context learning in conjunction with LLMs. This approach facilitates a deeper understanding of the relationships between chemical structures and their corresponding textual descriptions. Furthermore, ChemGPT [36] represents a variant of the GPT model specifically trained on chemical data. Through well-designed instructions, ChemGPT is capable of generating novel chemical structures and accurately predicting their properties. MolT5 [10] shows that LLMs can perform the cross-modal transition task between molecule and text (i.e. molecule captioning task and text-based molecule generation task), which is one of the most closely related attempts to ours. Besides, it should be noted that MolT5 still needs to be pre-trained on domain corpus and further fine-tuned for translating between molecules and natural language, leading to huge computational costs and harsh data requirements.
@@ 3 MolReGPT
Figure 2: The workflow of MolReGPT. MolReGPT consists of four main stages. In stage 1, Molecule-Caption Retrieval is employed to find $n$ best-matched examples from the local database. Then, in stage 2, Prompt Management helps construct the system prompt with the retrieved molecule-caption pairs. Following this, LLMs perform In-Context Few-Shot Molecule Learning based on the provided system prompt and user input prompt. Finally, Generation Calibration ensures the desired output.
Due to the huge computation and domain data labelling costs, training or fine-tuning LLMs on the domain-specific corpus in molecule discovery is often infeasible in practice. To address such limitations, we investigate leveraging the great capabilities of LLMs without changing the LLMs, where we propose a novel framework (MolReGPT) to equip ChatGPT with the ability of molecule-caption translation for molecule discovery. Specifically, in order to improve the quality of prediction, an In-Context Few-Shot Molecule Learning paradigm is introduced to teach ChatGPT to learn the molecule-caption translation task from context examples. The framework of MolReGPT is shown in Figure 2, consisting of four main stages: Molecule-Caption Retrieval, Prompt Management, In-Context Few-Shot Molecule Learning, and Generation Calibration, following the workflow of pre-processing, querying, and post-processing.